OpenStudy (anonymous):

Differentiate: y=cscxln(x^6/2x-1) I know the answer is y'=-cscxcotxln (x^6/2x-1) + cscx (6/x - 2/2x-1) I'm not sure how the last set of parenthesis came about, I thought the derivative would be 6x^5.

OpenStudy (loser66):

I answer you the last part only, is it ok?

OpenStudy (loser66):

that is ln (x^6 / 2x-1) = ln x^6 - ln (2x -1) . when you take derivative, \[(ln x^6)' =\]\[\frac{ 1 }{ x^6 }(6x^5)=\frac{ 1 }{ x }\] the same with the second part. so that's why you have it. :)

OpenStudy (anonymous):

Then why in the final answer is it (6/x)-(2/2x-1) ?

OpenStudy (loser66):

ok, the second part is (ln (2x-1))' = \[\frac{ 1 }{ 2x-1 }(2x-1)' = \frac{ 2 }{ 2x -1 }\] now combine them, ( let ....... csc x* aside, just calculate the last part is (ln (x^6/ 2x -1) )' as I calculate above, the derivative of that ln = (6/x) -(2/2x-1) . just write that answer after cscx* (.....) you have yours got it?

OpenStudy (anonymous):

Ahh gotcha thanks :)

OpenStudy (loser66):

ok, you must be something, I give out the shortcut and you got it quickly. :)

OpenStudy (anonymous):

I figured out the first part my brain just fizzled at the end. But thank you :) It's finals week lol